OpenStudy (anonymous):

Ok, I am working on circles in college algebra here. I am supposed to write this in standard form and determine both the center and radius if it exists. x^2+y^2-2x+6y-15=0

OpenStudy (anonymous):

Rewrite: x^2-2x+y^2+6y=15 Then complete the square: x^2-2x+1+y^2+6y+9=15+1+9 (x-1)^2+(y+3)^2=25

OpenStudy (anonymous):

(x - 1)^2 + ( y + 3)^2 = 5^2

OpenStudy (anonymous):

do you know how to "complete the square?"

OpenStudy (anonymous):

( 1, -3) are centre 's coordinates ........ 5 is the radius

OpenStudy (anonymous):

both answer are correct. do you know how to get them?

OpenStudy (anonymous):

@satellite... I thought I did but apparently I am not clear on it...

OpenStudy (anonymous):

lets rewrite as \[x^2-2x+y^2+6y=15\] as step one

OpenStudy (anonymous):

step two is to think "half of -2 is -1, and -1 squared is 1" half of 6 is 3 and 3 squared is 9

OpenStudy (anonymous):

yeah, I had it worked out that far...

OpenStudy (anonymous):

and then write \[(x-1)^2+(y+3)^2=15+1+9\]

OpenStudy (anonymous):

that is it. so you have the form you want it in.

OpenStudy (anonymous):

oh geez.... I was soooo stinkin' close, thanks!

OpenStudy (anonymous):

yw